OpenStudy (anonymous):

Solve the problem. Be sure to show all of your work. The wavelength, w, in meters per cycle, of a musical note is given by where r is the speed of sound in meters per second and f is the frequency in cycles per second. The speed of sound in air is 344 m/sec. What is the wavelength of a note whose frequency in air is 45 cycles per second? Round to the nearest tenth of a meter per cycle.

OpenStudy (amistre64):

you really should proof your questions ..... this says nothing

OpenStudy (anonymous):

?

OpenStudy (amistre64):

there are some sites that a copy paste from deletes vital stuff ..... like that 'given by' equation lol

OpenStudy (anonymous):

w=344/45

OpenStudy (anonymous):

i dont understand aything ur saying if u cat help dont eve bother

OpenStudy (anonymous):

thats the problem

OpenStudy (amistre64):

try to comprehend what im saying please ..... 'The wavelength, w, in meters per cycle, of a musical note is given by___equation is missing because you dint proof it____ where r is the speed....'

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

that part that tells you how to solve is ' IS NOT THERE'

OpenStudy (anonymous):

w= r/f

OpenStudy (amistre64):

thats better lol .....

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

amistre when you are done helping her can you check out my math problem

OpenStudy (amistre64):

i can try :)

OpenStudy (amistre64):

so its simply 344/45 right? = 7.6444444......

OpenStudy (anonymous):

yes and the tenth meter?

OpenStudy (anonymous):

=8?

OpenStudy (amistre64):

no, just 7.6 is what they want

OpenStudy (anonymous):

so where its say round i dont have to?

OpenStudy (amistre64):

you round it to the: 0.x ^ that spot

OpenStudy (amistre64):

7.64444 rounds to 7.6

OpenStudy (anonymous):

oh ok so its after the decimal point which cant be round cause the next number is lower than 5?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

k thaku